OpenStudy (anonymous):

WILL GIVE MEDAL!! If (7^2)p = 7^6, what is the value of p? 2 3 4 8

OpenStudy (campbell_st):

use the power of a power law for indexes \[(x^a)^b = x^{a \times b}\] so on the left side, you multiply the powers then you can equate them and solve for p

zepdrix (zepdrix):

Woops! I think he wants the addition rule, ya? Unless the problem was pasted incorrectly, it looks like this: \(\large\rm (7^2)p = 7^6\)

OpenStudy (anonymous):

yes that's right @zepdrix

zepdrix (zepdrix):

Oh maybe Camp takin' a nap :) I'll show you the rule in case you forgot.

OpenStudy (campbell_st):

well I read the problem as \[(7^2)^p = 7^6\]

zepdrix (zepdrix):

Exponent Addition Rule:\[\large\rm x^{a}\cdot \color{orangered}{x^{b}}=x^{a+b}\]Do you see how we can maybe apply this to your problem? :o \[\large\rm (7^2)\color{orangered}{p} = 7^6\]

zepdrix (zepdrix):

oh wait, you're right camp! look at the multiple choice. They would have to be powers of 7 to be addition rule

OpenStudy (campbell_st):

so then the question should be \[(7^2) \times 7^p = 7^6\]

zepdrix (zepdrix):

hehe, my bad :D

OpenStudy (campbell_st):

if you use what @zepdrix originally posted you would be solving \[p = 7^4\]

OpenStudy (anonymous):

no @campbell_st @zepdrix it suppose to be (7^2)p = 7^2 the p isn't an exponent it stays the way I typed it

zepdrix (zepdrix):

Hmm, that doesn't quite make sense with your choices though :( Thinking...

OpenStudy (anonymous):

that's what the question is @zepdrix

OpenStudy (campbell_st):

so the obvious choice is \[(7^2)^p = 7^6\] or possibly \[(7^2) \times (7^p) = 7^6\]

OpenStudy (campbell_st):

o well, good luck with the answer, the way you posted the question \[(7^2)p = 7^6\] doesn't have a solution in your choices

OpenStudy (anonymous):

no @campbell_st the p is NOT an exponent. nevermind I already got the answer. its 3

zepdrix (zepdrix):

If it's 3, then your question was: (7^2)^p. p being an exponent. must've been a typo or something on the website/book that you're looking at. \[\large\rm (7^2)^3=7^{2\cdot3}=7^6\]